A sample of krypton gas occupies a volume of 9.84 L at 60 °C and 0.5360 atm.
If the volume of the gas sample is increased to 10900 mL, while its temperature is increased to 133 °C, the resulting gas pressure, in torr, will be_____ torr.
1
Expert's answer
2020-04-28T12:21:08-0400

(P1V1)/T1= (P2V2)/T2

P1=0.5360 atm

V1=9.84 lit

T1 = (60+273) K

= 333 K

V2= 10.9 lit

T2 = (133+273) K

=406 K

P2 =( 0.560*9.84*406)/(333*10.9) atm

=0.616 atm

=468 torr (1atm= 760 torr)